Find the missing coordinates of the given points
5x - 7y = 35
(0,),(0,)

Answer by stanbon(75887)

Find the missing coordinates of the given points
5x - 7y = 35
(0,)
If x = 0, -7y = 35 so y = -5
-------------------------------------
(,0)
If y = 0, 5x = 35 so x = 7
========================================
Cheers,
Stan H.